Solve by Addition/Elimination Method:4x+8y+8=0
6y=3x+8

Answer by mananth(16946)

4 x + 8 y = -8 .............1

-3 x + 6 y = 8 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by -4
12 x 24 y = -24
12 x -24 y = -32
Add the two equations
24 x = -56
/ 24
x = -7/3
plug value of x in (1)
4 x + 8 y = -8
-9 + 8 y = -8
8 y = -8 + 9
8 y = 1
y = 1/6
x=-7/3 y =1/6